OpenStudy (anonymous):

what is the square root of 108b^4???!!! Help simplify the radical expression! PLEASE?!

OpenStudy (anonymous):

\[\sqrt{108b ^{4}}\]

OpenStudy (anonymous):

\[6b \sqrt{3}\]???

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{ 108b^{4} }\] \[\Large \sqrt{ 36*3*b^2*b^2 }\] \[\Large \sqrt{ 36 }* \sqrt{ 3 }* \sqrt{ b^2 }* \sqrt{ b^2 }\] \[\Large 6* \sqrt{ 3 }* b * b\] \[\Large 6* \sqrt{ 3 }* b^2\] \[\Large 6b^2\sqrt{ 3 }\] So \[\Large \sqrt{ 108b^{4} }\] simplifies to \[\Large 6b^2\sqrt{ 3 }\]

jimthompson5910 (jim_thompson5910):

You are very close, the b in the answer should be squared

OpenStudy (anonymous):

oops could u check \[5\sqrt{320}\] to make sure im doing this right?

jimthompson5910 (jim_thompson5910):

\[\Large 5\sqrt{320}\] \[\Large 5\sqrt{64*5}\] \[\Large 5\sqrt{64}*\sqrt{5}\] \[\Large 5*8*\sqrt{5}\] \[\Large 40\sqrt{5}\] So \[\Large 5\sqrt{320}=40\sqrt{5}\]

OpenStudy (anonymous):

thats what i got! one more? \[3\sqrt{12x ^{2}}\]

jimthompson5910 (jim_thompson5910):

\[\Large 3\sqrt{12x^2}\] \[\Large 3\sqrt{4*3*x^2}\] \[\Large 3\sqrt{4}*\sqrt{3}*\sqrt{x^2}\] \[\Large 3*2*\sqrt{3}*x\] \[\Large 6*\sqrt{3}*x\] \[\Large 6x\sqrt{3}\] So \[\Large 3\sqrt{12x^2}\] simplifies to \[\Large 6x\sqrt{3}\]

OpenStudy (anonymous):

thank you so much for your help and if thatquestion had been \[-3\sqrt{12x ^{2}}\]would the answer be \[-6x \sqrt{3}\]?

jimthompson5910 (jim_thompson5910):

yes that's exactly right

OpenStudy (anonymous):

thank you so much i grasp the concept well now(:

jimthompson5910 (jim_thompson5910):

that's great

jimthompson5910 (jim_thompson5910):

you're welcome